1/1

Manfred Hausmann,

Maple Bridge Night Mooring

übersetzt von Zhao Yanchun

The moon setting, crows cawing, frost overfilling the skies, The maples riv’rside, lamps aboard, sad the fishermen’s eyes. The toll from the Cold Mountain Temple outside of Suchow Reaches my boat at midnight and floods my sleepless sorrow.

Übersetzung des deutschen Gedichtes “An der Ahornbrücke” von Manfred Hausmann (∗1898-09-10 – †1986-08-06) ins Englische durch Zhao Yanchun.

Arg-911-1824 (2013-07-27 10:23:29)

http://www.poezio.net/version?poem-id=911&version-id=1824